OpenStudy (anonymous):

Salina was at a party and she noticed that the room had 54 balloons. For every red balloon there were 2 yellow balloons and 3 blue balloons. How many blue balloons were there?

OpenStudy (anonymous):

how do i figure this when it doesnt tell me how many red balloons there are

OpenStudy (anonymous):

your answer is A

OpenStudy (swissgirl):

well for starters we know that there are 54 balloons

OpenStudy (swissgirl):

ok now as you said we dont know how many red balloons we have so lets say we have x red balloons

OpenStudy (anonymous):

and that the answer is A:)

OpenStudy (swissgirl):

Now for every 1 red balloon we have 2 yellow balloons so if we have 2 red balloons we will have 2*2=4 yellow balloons. So basically there are twice the amount of red ballons yellow balloons =2x

OpenStudy (anonymous):

kelly stop texting stupid stuff that doesnt make sense

OpenStudy (swissgirl):

Now for every red balloon there are 3 blue balloons blue balloons=3x Now in total we will have x+2x+3x=54

OpenStudy (swissgirl):

Now can you solve for x to find out how many red balloons there are

OpenStudy (swissgirl):

so what did you get for x?

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

x=9 right

OpenStudy (swissgirl):

yup now we know that there are 9 red balloons

OpenStudy (swissgirl):

so how many blue balloons would there be?

OpenStudy (anonymous):

27 :):):)

OpenStudy (swissgirl):

YAHHHOOOOOOOOO

OpenStudy (anonymous):

At an amusement park, each premium ride costs $12 and each gold star ride is $8. Stan spends $96 on tickets and he gets 3 more premium rides than gold star rides. How much does he spend on gold star rides?

OpenStudy (swissgirl):

Lets see what @robtobey has to say. Maybe i did something incorrectly

OpenStudy (anonymous):

27 blue balloons. There are 54/6=9 groups of red/yellow/blue balloons and in every balloon group there are 3 blue balloons. 9*3=27

OpenStudy (swissgirl):

ohhh yay we got the correct answer

OpenStudy (anonymous):

i got 24$

OpenStudy (anonymous):

on the last question i just asked can someone make sure i did that right please

OpenStudy (anonymous):

At an amusement park, each premium ride costs $12 and each gold star ride is $8. Stan spends $96 on tickets and he gets 3 more premium rides than gold star rides. How much does he spend on gold star rides?

OpenStudy (anonymous):

$24

OpenStudy (anonymous):

??????

OpenStudy (swissgirl):

yes I am here

OpenStudy (swissgirl):

Ok lets set up our equation first

OpenStudy (swissgirl):

ok so we purchase $96 worth of tickets we buy x amount of gold tickets and x+3 premium tickets ok

OpenStudy (swissgirl):

now how would you set up our equation?

OpenStudy (swissgirl):

did i lose you?

OpenStudy (anonymous):

no i just already figured out that i was right

OpenStudy (anonymous):

lol

OpenStudy (swissgirl):

yes you are correct sorryyyyyy

OpenStudy (swissgirl):

I thought premium rides hahahahaahah

OpenStudy (swissgirl):

ok so you got it :) U R AWESOME

OpenStudy (anonymous):

Let gsr be the number of gold start rides that Stan bought. Solve the following for gsr: 12 (gsr + 3) + 8 gsr = 96 gsr = 3 8*3 = $24

OpenStudy (anonymous):

Miguel is making smoothies out of yogurt and juice to serve to his friends, and he needs to make 12 cups. The ratio of yogurt to juice is 2 cups to 1 cup. How much yogurt will he use?

OpenStudy (swissgirl):

ok so lets label juice as x and so for every cup of juice we have 2 cups of yogurt Yogurt=2x How would we set up our equation?

OpenStudy (anonymous):

2 times 12 right but the only thing about that is that these are the choices i have for the answer 3,4,8,9

OpenStudy (swissgirl):

x+2x=12 3x=12 x=4 So he uses 4 glasses of juice Yogurt is 2x=2*4=8 So he will use 8 cups of yogurt

OpenStudy (anonymous):

oh ok that makes more sense than what i tried doing lol

OpenStudy (swissgirl):

lol

OpenStudy (anonymous):

Latonya bought flowers for her mother. She spent $45. Each tulip was $1 and each rose was $2. She bought 3 more tulips than roses. Let t be the number of tulips. Which algebraic sentence would help you figure out how many tulips Latonya bought?

OpenStudy (anonymous):

A. t + 2(t + 3) = 45 B. t + 2(t – 3) = 45 C. 2t (t + 3) = 45 D. -3t + 2t = 45

OpenStudy (swissgirl):

Ok lets skip the algebraic sentence first and try to set up the equation

OpenStudy (swissgirl):

ohh they provided that answer

OpenStudy (swissgirl):

ok so you tell me.

OpenStudy (anonymous):

one sec let me try to figure that out haha in my HEAD lol

OpenStudy (swissgirl):

ok there is something tricky here. Should I give you a hint

OpenStudy (anonymous):

YES PLEASe

OpenStudy (anonymous):

i think it would either be A or C is that correct at least?

OpenStudy (swissgirl):

ok t=tulips we have 3 more tulips than roses Now that is another way of saying that we have 3 less roses than tulips

OpenStudy (swissgirl):

Sorry thats incorrect :(

OpenStudy (swissgirl):

ok lets work through it together

OpenStudy (swissgirl):

ok we have t amount of tulips and how many roses do we hav

OpenStudy (anonymous):

2?

OpenStudy (swissgirl):

well they cost $2 per rose

OpenStudy (swissgirl):

but how many roses do we have in comparison to tulips

OpenStudy (swissgirl):

if we have t tulips then how many roses do we have Do we have more roses than tulips Or less roses than tulip

OpenStudy (swissgirl):

She bought 3 more tulips than roses. meaning that we have more tulips than roses correct?

OpenStudy (swissgirl):

if we have t amount of tulips we have 3 less roses

OpenStudy (anonymous):

$28 on roses and $17 on tulips right

OpenStudy (anonymous):

A. t + 2(t + 3) = 45 B. t + 2(t – 3) = 45 C. 2t (t + 3) = 45 D. -3t + 2t = 45 i still think the answer is either a or c but do those make sense

OpenStudy (swissgirl):

Ok you got the correct answer so what was your equation

OpenStudy (swissgirl):

How did you get your answer?

OpenStudy (anonymous):

i didnt write down any equation but since i am adding so many ts more than roses i did the t+3 obviously and i would assume t+3 at the beginning because its a dollar plus two one for the tulip and two for the rose so the equations just look like a or c could work for solving it am i right?

OpenStudy (swissgirl):

ok let me show what I did We have t tulips and t-3 roses

OpenStudy (anonymous):

i added 3 together until i got to 42 and then added just the three tulips to get 45

OpenStudy (swissgirl):

Since we have 3 roses less than tulips

OpenStudy (swissgirl):

so we have t+2(t-3)=45 and since we have 17 tulips plug in t=17 and you will see you will get the correct answer

OpenStudy (swissgirl):

so our answer is B

OpenStudy (swissgirl):

this one was confusing We had t tulips and we didnt have more tulips than roses but we had 3 less roses than tulips Roses=t-3

OpenStudy (anonymous):

A little late for the yogurt/juice problem. Let y and j be then number of cups of yogurt and juice, respectively, to make 12 cups of smoothies. Solve the following for y and j. The dimensions for the first equation is cups and the second equation is dimensionless, a pure ratio.\[\left\{y+j=12,\frac{y}{j}=2\right\}\]\[\{y=8 \text{ cups},j=4 \text{ cups}\} \]

OpenStudy (anonymous):

oh ok so thats where i got confused lol i was assuming it was +3 but -3 makes more sense

OpenStudy (swissgirl):

lol it was a little tricky

OpenStudy (anonymous):

thankss for everything LATER ;]

OpenStudy (swissgirl):

Bye :)